Jill purchased a share one year ago for $13.19. The share is now worth $14.81, and the total return to Jill for owning the share was 23.6 per cent. The dollar amount of dividends that she received for owning the share during the year is (expressed in dollars to the nearest cent; don't use $sign or commas eg 50 cents is 0.50)
Answer #1

Initial Price = $ 13.19 and Final Price = $ 14.81

Capital Gains in % = [(14.81-13.19) / 13.19] x 100 = 12.282 %

Total Return = Capital Gains Return + Dividend Return = 23.6 %

Dividend Return = 23.6 - 12.282 = 11.318 %

$ Amount of Dividend = 0.11318 x 13.19 = $ 1.493 ~ $ 1.49